If 220 ml of 1 F AgNO2 is required to react quantitatively with 150 ml of KCN solution according to the reaction expressed by the equation Ag+ + 2CN- = Ag(CN)2-


1
Expert's answer
2021-03-03T01:31:27-0500

AgNO2 + 2KCN = (Ag(CN)2)K+ KNO2

n (AgNO2) = 2 x n (KCN)

n = CM x V

1 x 0.220 = 2 x (CM(KCN) x 0.150)

CM(KCN) = (1 x 0.220)/ (2 x 0.150) = 0.733 M